Excess concentrated hydrochloric acid was added to 2.0 of potassium tetra oxo manganette (vii) and chlorine gas was evolved and collected as stp the equation is below
2kmno4+16Hcl =2 kcl+ 2mncl2 +8h2O+ 5 cl2
1
Expert's answer
2020-01-22T04:18:32-0500

n(KMnO4)=2.0/158,03=0.01265 mol

n(Cl2)=2.5(0.01265)=0.031625 mol

V(Cl2)=22.4(0.031625)=0.7084 l

Answer: V(Cl2)=0.7084 l